### v1

# Multi-type branching processes with immigration generated by point processes

Abstract
Following the pivotal work of Sevastyanov [38], who considered branching processes with homogeneous Poisson immigration, much has been done to understand the behaviour of such processes under different types of branching and immigration mechanisms.
Recently, the case where the times of immigration are generated by a non-homogeneous Poisson process was considered in depth. In this work, we try to demonstrate how one can use the framework of point processes in order to go beyond the Poisson process. As an illustration, we show how to transfer techniques from the case of Poisson immigration to the case where it is spanned by a determinantal point process.
Keywords: Branching processes with immigration, Point processes.
MSC2020 Classification: 60J80, 60G55
1 Introduction
The main purpose of this paper is to examine the impact of immigration on behaviour of the continuous-time Markov branching processes, both single- and multi-type, focussing on immigration types that differ from the Poisson model. Although classical Poisson processes are the most commonly used to model the immigration component, they have certain limitations in practical applications. These include the assumption of equidispersion, which may not always hold, and their time-homogeneity, which is addressed by introducing time-non-homogeneous intensity rates. Recently, such branching models with immigration at the jump times of a non-homogeneous Poisson process were considered in depth by Mitov et al. [17] in the critical case and later on by Slavtchova-Bojkova et al. [42, 43] in the non-critical cases. In these studies, the efforts were concentrated mainly on analysing the asymptotic behaviour of the processes for various rates of the Poisson measure, under the assumption that these rates are asymptotically equivalent to either exponential or regularly varying functions. Consequently, results similar to the strong law of large numbers and central limit theorems were established.
Multi-type Markov branching processes were first introduced by Kolmogorov and Dmitriev [20], marking the term branching process as one of the earliest concepts in the literature. The notion of branching processes with immigration was later formalised by Sevastyanov [38], who explored a single-type Markov process with immigration driven by a time-homogeneous Poisson process, and derived limiting distributions for subcritical, supercritical, and critical cases.
Since then, numerous extensions of branching processes with immigration have been developed and thoroughly explored. Seminal reviews by Sevastyanov [39], and Vatutin and Zubkov [45, 46], have highlighted many key results. More recent advancements have been made by Barczy et al. [5, 6], González et al. [13], and Li et al. [24], among others, who have expanded the theory of these processes.
Branching processes with time-non-homogeneous immigration were first introduced by Durham [10] and Foster and Williamson [12]. Comprehensive reviews of these models can be found in the monographs by Badalbaev et al. [4] and Rahimov [36], as well as in Rahimov’s more recent review paper [37]. In the recent work of Rabehasaina and Woo [35], the model described in Mitov et al. [17] and Slavtchova-Bojkova et al. [42, 43] were considered and the established limit results were acquired by means of characteristic functions.
In this work, we aim to introduce a general framework for the Laplace transforms of random point processes within the context of branching processes. We believe that this approach will be valuable for researchers looking to extend their work beyond the Poisson process. For example, our method simplifies the derivation of the generating function for a branching process with immigration, which often serves as the foundation for many studies. Traditionally, this derivation was based on the exact distribution of points over a given interval, which is tractable in the case of a Poisson process, as demonstrated in [17, Theorem 1], [35, Lemma 3.1], and [7, Theorem 1]. Our framework provides a more unified and often simpler approach to addressing problems involving immigration, as one can see from the proof of the third item of Theorem 4.5 and similar results for Poisson immigration, such as [17, Theorem 8] and [35, Theorem 3]. Due to the tractability of the framework, we correct minor errors from the last two referenced results, see Remark 4.6.
The paper is structured as follows: in Section 2 we define the framework of point processes, and give a few examples of such objects, and in Section 3 we introduce single- and multi-type branching processes with immigration. Section 4 contains results for the probability generating functions of the branching process with immigration spanned by a general point process, and asymptotic results in the case where the immigration is generated by a Determinantal Point Process (DPP). More precisely, in Theorem 4.1 the generating functions of branching processes with immigration spanned by DPPs, Cox process and Fractional Poisson Process (FPP) are derived in the single-type case, while Theorem 4.2 presents the multivariate analogue. In Subsection 4.3, for the case of DPPs immigration, we establish the equations for the mean and covariances of the process. Furthermore, in Theorem 4.4, we prove limit results analogous to those in the case without immigration, covering sub-, super-, and critical cases, respectively. Section 5 contains the proofs of the new results.
2 Point processes on the real line
As we will be interested in point processes (PPs) describing the times of immigrants joining a branching process, we define the basic notions for PPs on $\mathbb{R}$ (or just PPs from now on) although the formalism for defining them on an arbitrary Polish space is similar.
Informally, we can consider the point process as a random collection of points; however, to formalise this idea, the basic construction is via the so-called random measures.
We follow the presentation of [22, Chapter 2.1], however, another suitable reference is the excellent monograph by Baccelli et al. [3].
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\mathcal{B}(\mathbb{R})$ be the usual Borel $\sigma$-algebra on $\mathbb{R}$, $\mathbf{N}_{<\infty}$ be the set of measures on $\mathbb{R}$ such that for each $B\in\mathcal{B}(\mathbb{R}),\Phi(B)\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$, and $\mathbf{N}$ the set of measures which can be represented as a countable sum of elements of $\mathbf{N}_{<\infty}$. Let $\mathcal{N}$ be the $\sigma$-algebra generated by the sets
$\{\Phi\in\mathbf{N}:\Phi(B)=k\text{ for some }B\in\mathcal{B}\text{ and }k\in\mathbb{N}_{0}\}.$
We call $\Phi$ a random measure or a PP if it is a random element of $(\mathbf{N},\mathcal{N})$, that is, a measurable mapping $\Phi:\Omega\to\mathbf{N}$. In this work, we will consider PPs which are proper, i.e., such that there exist random variables $\kappa,X_{1},X_{2},\dots$ such that
$\Phi=\sum_{i\leq\kappa}\delta_{X_{i}},$
where $\delta_{x}$ is the Dirac mass at $x$. For a deeper and more general presentation, see [22, Chapter 2.1], [3, Chapter 1] or [18, Chapter 2].
Some of the characteristics which help describing a PP include:
  1. i)
    its intensity measure $\Lambda$, defined by $\Lambda(B)=\mathbb{E}\left[\Phi(B)\right]$, where $B\in\mathcal{B}(\mathbb{R})$;
  2. ii)
    its Laplace functional, which characterises the process completely, for all test functions $f$,
    $\mathcal{L}_{\Phi}(f):=\mathbb{E}\left[e^{-\int f\mathrm{d}\Phi}\right]=\mathbb{E}\left[e^{-\sum_{i}f(X_{i})}\right].$ (2.1)
    The set of test functions can be all positive measurable ones, like in the case of a Poisson process, or these of compact support, like in the case of determinantal point processes;
  3. iii)
    its joint intensities $\rho_{k}$, if they exist, defined as $\rho_{k}:\mathbb{R}_{+}^{n}\to\mathbb{R}_{+}$ such that for each disjoint $B_{1},\dots,B_{k}\in\mathcal{B}(\mathbb{R})$,
    $\mathbb{E}\left[\prod_{i=1}^{k}\Phi(B_{i})\right]=\int_{B_{1}\times\dots\times B_{n}}\rho_{k}(x_{1},\dots,x_{k})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{k}).$
2.1 Poisson processes
Probably the most used point process, due to its mathematical tractability, is the Poisson one, which is characterised by the property that its intensity measure $\Lambda$ is such that:
  1. i)
    for every $B\in\mathcal{B}(\mathbb{R})$, $\Phi(B)\sim Pois(\Lambda(B))$ and
  2. ii)
    for every disjoint $B_{1},\dots,B_{m}\in\mathcal{B}(\mathbb{R})$, $\Phi(B_{1}),\dots,\Phi(B_{n})$ are independent.
It is a consequence that the Laplace transform of this PP is then
$\mathcal{L}_{\Phi}(f)=\exp\left(-\int_{\mathbb{R}}\left(1-e^{-f(x)}\right)\Lambda(\mathrm{d}x)\right).$
In the following three sections, we list some PPs which are often used and cover, respectively, repulsive, clustering, and heavy-tailed behaviour of the interarrival times.
2.2 Determinantal point processes (DPPs)
DPPs were introduced by Macchi [25] under the name fermion process due to their repulsive behaviour. Since then, they appear in various contexts as describing fundamental objects in random matrix theory, random analytic functions, statistical mechanics, and machine learning [21]. For an introduction to their formalism, we refer to [3, Chapter 5], and for an overview, to the survey [8].
We will call a point process (in this work mainly on $\mathbb{R}_{+}$) $(\Lambda,K)$-determinantal with a directing measure $\Lambda$ if its joint densities satisfy, for $x_{i}\in\mathbb{R}$,
$\rho_{k}(x_{1},\dots,x_{k})=\det(K(x_{i},x_{j}))_{1\leq i,j\leq k}$
with $K$ a Hermitian, locally trace class, integral operator on $L^{2}(\mathbb{R},\Lambda)$. Its Laplace transform is given, for a function $f:\mathbb{R}_{+}\to\mathbb{R}$ with bounded support and
$\varphi_{n}(x_{1},\dots,x_{n}):=\prod_{i=1}^{n}(1-e^{-f(x_{i}))}),$
by
$\mathcal{L}_{\Phi}(f)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{\mathbb{R}^{n}}\varphi_{n}(x_{1},\dots,x_{n})\rho_{n}(x_{1},\dots,x_{n})\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}),$ (2.2)
see [3, Proposition 5.1.18, Corollary 5.1.19]. In particular, note that if we choose the kernel $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, then $\rho_{n}\equiv 1$ for all $n$, and we obtain the usual $\Lambda$-Poisson process.
2.3 Cox Processes
Cox processes, also known as doubly stochastic Poisson processes, are commonly used for modelling phenomena where event occurrence is influenced by underlying random factors, as demonstrated in [31, 32, 48] and more recently in [9], which discuss their applications in spatial and spatio-temporal data analysis. These processes exhibit behaviour that contrasts with the repulsive nature of determinantal point processes (DPPs), making them particularly suitable for capturing patterns where events tend to cluster.
Formally, a Cox process is constructed by first drawing an intensity measure $\Lambda$ from a random distribution $\eta$ defined on the space of $\sigma$-finite measures on $\mathbb{R}$. Given this intensity measure, a Poisson process with intensity $\Lambda$ is then generated. Formally, we can describe this construction as:
$\Lambda$ $\sim\eta,$
[…]
2.4 Fractional Poisson processes
The fractional Poisson process (FPP) was introduced by Mainardi et al. [26] as a generalisation of the standard homogeneous Poisson process, which is non-Markovian and has heavy-tailed interarrival times; see also the monograph by Meerschaert and Sikorskii [28]. This process is of renewal type, that is, we can consider that the atoms of the associated random measure are situated at $X_{n}=T_{1}+\dots+T_{n}$ with iid $T_{i}$, such that $\mathbb{P}(T_{n}>t)=E_{\beta}(-\lambda t^{\beta})$ for $\beta\in(0,1)$ and
$E_{\beta}(z):=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(1+\beta k)}$
denote the Mittag-Leffler function. It should be noted that $\mathbb{E}\left[T_{i}\right]=\infty$ and that by the work of Meerschaert et al. [27], the counting process $N_{t}:=\Phi((0,t])$ can be constructed as time-changed $\Lambda$-Poisson process, for $t\geq 0$ and $\alpha\in(0,1)$,
$N_{\alpha}(t)=N(Y_{\alpha}(t)),$ (2.3)
where $N$ is the classical homogeneous Poisson process with parameter $\lambda>0$, and $Y_{\alpha}$ is the inverse of an independent $\alpha$-stable subordinator.
3 Branching processes with immigration
In this section we present the branching process with immigration and fix the relevant notation.
3.1 Single-type processes
We denote by $Z(t)$ the number of particles at time $t$ and outline that we use $G_{X}$ for the generating function of a random variable $X$.
The underlying branching process without immigration $Z_{\times}$ will be a Markov branching process in continuous time with starting state $Z_{\times}(0)=1$. Therefore we can construct the process with immigration $Z$ as follows:
  • At times $T_{1}<T_{2}<\dots$, new particles (immigrants) $I_{1},I_{2},\dots$ arrive with $I_{k}\sim I$ iid. We assume that $\mathbb{E}\left[I^{2}\right]$ is finite.
  • Each particle evolves independently of the others and at $Exp(1/\mu)$ time dies and produces $\nu$ new particles. It is then known, see [2, p.102], that
    $\frac{\partial}{\partial t}G_{Z_{\times}}(t,s)=f_{\nu}(G_{Z_{\times}}(t,s)),$
    where $G_{Z_{\times}}(t,s):=G_{Z_{\times}(t)}(s)$ and $f_{\nu}(s)=(G_{\nu}(s)-s)/\mu$. Also, $M_{\times}(t):=\mathbb{E}\left[Z_{\times}(t)\right]=e^{\rho t}$ with $\rho:=\mu(\mathbb{E}\left[\nu\right]-1)$. The process is called subcritical, critical or supercritical according to whether $\mathbb{E}\left[\nu\right]$ is, respectively, less, equal, or larger than 1.
Therefore, the process with immigration can be represented as
$Z(t)=\sum_{i:T_{i}\leq t}{Z}^{(i)}_{\times}(t-T_{i}),$
where ${Z}^{(i)}_{\times}$ are iid copies of the process without immigration $Z_{\times}$ started with ${Z}^{(i)}_{\times}(0)=I_{i}$.
3.2 Multi-type processes
Consider now a generalisation of the process presented above which is composed of $d$ types of particles. We note that we will use bold symbols, $\boldsymbol{\nu}$ for a vector, and, unless stated otherwise explicitly, $\nu_{i}$ will denote the value of its $i$th coordinate. We define the generating function in this case as
$G_{\boldsymbol{\nu}}(\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}^{d}}\mathbb{P}\left(\boldsymbol{\nu}=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}}.$
The multi-type branching process $\boldsymbol{Z}$ with $\boldsymbol{Z}(0)=\boldsymbol{I}^{(0)}\sim\boldsymbol{I}$, is described by:
  • At times $0=T_{0}<T_{1}<T_{2}<\dots$, iid immigrants $\boldsymbol{I}^{(0)},\boldsymbol{I}^{(1)},\boldsymbol{I}^{(2)},\dots\sim\boldsymbol{I}$ join the population described by the branching process $\boldsymbol{Z}.$ We assume that $\mathbb{E}\left[I_{i}^{2}\right]$ are finite for all $i$.
  • Each particle of type $i$ lives $Exp(\mu_{i})$ time after which it dies and produces new ones with distribution $\boldsymbol{\nu}_{i}$ of their counts, that is to say $({\nu}_{i})_{j}$ of type $j$.
Assumptions: In this work we assume that
  1. (i)
    the immigrants and progeny, ${I_{i}}$ and ${(\nu_{i})}_{j}$ have finite second moments;
  2. (ii)
    the matrix $M:={(\nu_{i})}_{j}$ is such that for some $p\in\mathbb{N}$, $M^{p}$ has strictly positive terms;
  3. (iii)
    and ${(\nu_{i})}_{j}$ have non-lattice distributions.
Similarly to the first case,
$\boldsymbol{Z}(t)=\sum_{i:T_{i}\leq t}\boldsymbol{Z}^{(i)}_{\times}(t-T_{i}),$
where ${Z}^{(i)}_{\times}$ are iid processes with law $\boldsymbol{Z_{\times}}$, that is, without immigration and started with $\boldsymbol{I^{(i)}}$. This gives the generating function of the process, with $G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=G_{\boldsymbol{Z}(t)}(\boldsymbol{s})$ and similarly for $G_{\boldsymbol{Z}_{\times}}(t,\boldsymbol{s}$),
$G_{\boldsymbol{Z}}(t,\boldsymbol{s})=\mathbb{E}\left[\prod_{i:T_{i}\leq t}G_{\boldsymbol{Z}^{(i)}_{\times}}(t-T_{i},\boldsymbol{s})\right]=\mathbb{E}\left[\exp\left(\sum_{i:T_{i}\leq t}\ln\left(G_{\boldsymbol{Z}_{\times}}(t-T_{i},\boldsymbol{s})\right)\right)\right].$ (3.1)
Translating the last in the language of the Laplace transform of the random measure $\Phi$ governing the process $T_{i}$, substituting in (2.1), we get that
$G_{\boldsymbol{Z}}(t,\boldsymbol{s})=\mathcal{L}_{\Phi}(f_{t}),\quad\text{with}\quad f_{t}(x):=-\ln\left(G_{\boldsymbol{Z}_{\times}}\left(t-x,\boldsymbol{s}\right)\right)\mathbbm{1}_{\left\{x\leq t\right\}}.$
If we want to work instead with the Laplace transform,
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp(\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle)\right],$
[…]
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s})=\mathcal{L}_{\Phi}(g_{t}),\quad\text{with}\quad g_{t}(x):=-\ln\left(\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t-x,\boldsymbol{s}\right)\right)\mathbbm{1}_{\left\{x\leq t\right\}}.$ (3.2)
4 Asymptotic behaviour at infinity of the branching process with immigration
4.1 Single-type processes
The continuous-time Markov branching process without immigration is presented in Athreya and Ney [2, Chapter III]. We refer again to Rahimov [37] for an extensive recent review of branching processes with immigration and to Vatutin and Zubkov [45] for a review of the classical results until 1983.
A standard argument using martingale theory shows that there is a random variable $W$ such that for the process without immigration $Z_{\times}$, there exists a real-valued random variable $W_{\times}$ such that
$\frac{Z_{\times}(t)}{e^{\rho t}}\xrightarrow[t\to\infty]{a.s.}W_{\times}.$ (4.1)
In the supercritical case, it is natural to characterise some of the properties of $W_{\times}$, e.g., the existence of a density. In the sub- and critical cases, this $W_{\times}$ is a.s. equal to 00 and a common question is to analyse the so-called Yaglom limit as $t\to\infty$ of quantities such as $\mathbb{P}(Z(t)>f(t)|Z(t)>0)$ for some $f$, see [2, Chapter III, Theorems 2-4]. Similar results are also available for the case of age-dependent processes, see [2, Chapter IV]
When an immigration component is added, it is often possible to extract asymptotic results for $Z$ using the information for the underlying process $Z_{\times}$ through representation (3.1). The case of Poisson immigration is considered when the underlying process is subcritical in [16] , critical in [29, 30, 44], and supercritical in [15].
We provide the generating functions in the cases of DPPs, Cox, and FPPs. After choosing an exact model, this can provide the limit of a scaled $Z$ through the results for $Z_{\times}$. We also note that it is possible to work with the generating function of the process starting with exactly one particle $Z_{\times}^{(1)}$, that is $Z_{\times}^{(1)}(0)=1$, via the relation
$G_{Z_{\times}}(s)=G_{I}\left(G_{Z_{\times}^{(1)}}(s)\right).$ (4.2)
Theorem 4.1.
  1. 1.
    In the case of an immigration spanned by a $(\Lambda,k)$-DPP process, $\mathbb{E}\left[s^{Z(t)}\right]$ is equal to
    $1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\prod_{i=1}^{n}\left(1-G_{Z_{\times}}(t-x_{i},s)\right)\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}).$ (4.3)
  2. 2.
    In the case of an immigration spanned by a Cox process with a directing measure $\eta$,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}_{\eta}\left[\exp\left(-\int_{(0,t]}\left(1-G_{Z_{\times}}(t-x,s)\right)\Lambda(\mathrm{d}x)\right)\right]}.$
  3. 3.
    In the case of an immigration spanned by an $\alpha$-FPP,
    $\mathbb{E}\left[s^{Z(t)}\right]={\mathbb{E}_{\eta}\left[\exp\left(-\int_{(0,Y_{\alpha}(t)-]}\left(1-G_{Z_{\times}}\left(t-Y^{-1}_{\alpha}(t),s)\right)\right)\Lambda(\mathrm{d}t)\right)\right]}.$
Proof.
For the first item, we substitute the Laplace transform of DPPs (2.2) in (3.2). For the second, we use that Cox processes are Poisson ones with a random intensity. For the third item, we use the representation of FPPs as time-changed Poisson processes from (2.3). The last gives that if $\Phi$ is the Laplace transform of an $\alpha$-FPP,
$\mathcal{L}_{\Phi}\left(f\right)=\mathbb{E}\left[\exp\left(-\int\left(1-e^{-f\left(Y_{\alpha}^{-1}(t)\right)}\right)\right)\Lambda(\mathrm{d}t)\right].$
The object $Y_{\alpha}^{-1}(t)$ is the inverse stable subordinator and its analysis is often crucial in understanding the fractional Poisson process after the work [27] showed its relevance. Recently, the asymptotics of inverse subordinators were well understood under very mild assumptions by Ascione et al. in [1]. Substituting in (3.2), we obtain the third item of the theorem. ∎
To make a connection with the previously presented classes of point processes, the second item of the last theorem was proved by Butkovsky [7] and applied in the context of branching processes.
As for the fractional point process, since it is a renewal process, we can apply the results of Kaplan and Pakes [19, 33] for the sub- and supercritical cases, see also [34]. The only specificity one needs to take care of is that the interarrival times have infinite expectation, however the techniques are often the same, see [19, Remarks on page 379, 385, and 389].
4.2 Multi-type processes
[…]
$A_{ij}:=\mu_{i}^{-1}\left(\mathbb{E}\left[\left(\nu_{i}\right)_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}\right).$ (4.4)
We will suppose that $e^{A}$ is irreducible and let $\rho$ be its largest eigenvalue. The process is called sub-, super- or just critical according to $\rho$, respectively, less, larger or equal than 0. Let $\boldsymbol{u}$ and $\boldsymbol{v}$ be row vectors such that
$A\boldsymbol{u}^{t}=\rho\boldsymbol{u}^{t},\quad\boldsymbol{v}A=\rho\boldsymbol{v},\quad\sum u_{i}v_{i}=1,\quad\text{and}\quad\sum_{i}u_{i}=\\ 1,$
with $t$ denoting the transpose. Then, from [2, Theorem 1, p.205] or rewritten in the form [35, Lemma 1], there exists a real-valued random variable $W_{\times}$ such that
$\frac{\boldsymbol{Z}_{\times}(t)}{e^{\rho t}}\xrightarrow[t\to\infty]{a.s.}\boldsymbol{v}\cdot W_{\times},$ (4.5)
with $\cdot$ the usual multiplication, spelled out for better readability. We recall that $\boldsymbol{Z}_{\times}(t)$ is the process without immigration started with random number of particles according to the law $\boldsymbol{I}$. For completeness, we also note that the equivalent of the functional equation (4.2) is
$G_{\boldsymbol{Z}_{\times}}(s)=G_{\boldsymbol{I}}\left(G_{\boldsymbol{Z}^{(1)}_{\times}}(\boldsymbol{s}),\dots,G_{\boldsymbol{Z}^{(d)}_{\times}}(\boldsymbol{s})\right),$
with ${\boldsymbol{Z}^{(k)}_{\times}}$ a process without immigration started with exactly one particle of type $k$.
Furthermore, a convenient fact is that Theorem 4.1 holds when changing everything to its multidimensional equivalent (which we write in bold).
Theorem 4.2.
The multidimensional analogue of Theorem 4.1, which is obtained by substituting $s$ and $Z$ with, respectively, $\boldsymbol{s}$ and $\boldsymbol{Z}$, holds.
Therefore, in theory, we can use results for the process without immigration $\boldsymbol{Z}_{\times}$ to obtain information about $\boldsymbol{Z}$.
As noted previously, the case of immigration generated by a Poisson random measure is analysed by Yanev et al. in [17, 42, 43] and independently by Rabehasaina and Woo in [35]. For immigration times that form a renewal process, see [41]. The latter may be applied in the case of an immigration spanned by a FPP after checking that the used arguments can be modified to include the case of an infinite expectation of the inter-arrival times.
As an illustrative example of how ideas from the case of Poisson immigration can be modified to encompass the case of immigration spanned by a DPP, we generalise [35, Theorems 3, 4, and 5] (since the Poisson point process is a DPP), which are also available in [17, 42, 43].
4.3 Moments of the process
Theorem 4.3.
In the case of immigration spanned by a $(\Lambda,K)$-DPP, and denoting the $i$th coordinate with an index,
  1. (i)
    we have that
    $\mathbb{E}\left[{Z}_{i}(t)\right]=\int_{(0,t]}K(x,x)\mathbb{E}\left[Z_{\times,i}(t-x)\right]\Lambda(\mathrm{d}x),$
  2. (ii)
    and also
    $Cov(Z_{i}(t),Z_{j}(t))=\int_{(0,t]}K(x,x)$ $\mathbb{E}\left[Z_{\times,i}(t-x)Z_{\times,j}(t-x)\right]\Lambda(\mathrm{d}x)$
    $\!-\int_{(0,t]^{2}}K^{2}(x,y)\mathbb{E}\left[Z_{\times,i}(t-x)\right]\mathbb{E}\left[Z_{\times,j}(t-y)\right]\Lambda(\mathrm{d}x)\Lambda(\mathrm{d}y).$
Proof.
This follows by differentiating the Laplace transform of the process, derived as the multidimensional version of (4.3). We provide the complete calculation in Section 5.1.
Theorem 4.4.
In the case of immigration spanned by a $(\Lambda,K)$-DPP $\Phi$ with kernel $K$ such that $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)<\infty$, there exists an $\mathbb{R}^{d}$-valued random variable $\boldsymbol{W}$ such that
[…]
Moreover,
$\boldsymbol{W}\stackrel{{\scriptstyle d}}{{=}}\sum_{i}\boldsymbol{v}\cdot W^{(i)}_{\times}e^{-\rho T_{i}}$
with $W^{(i)}_{\times}$ iid copies of $W_{\times}$ and $T_{i}$ being the atoms of $\Phi$.
The proof of the last theorem is presented in Section 5.2.
For the next results, to introduce some regularity of the DPP, we assume it is stationary, meaning that for any $t\in\mathbb{R}$ and $\boldsymbol{x}\in\mathbb{R}^{n}$,
$D(\boldsymbol{x}+t)=D(\boldsymbol{x}),$ (4.7)
where $\boldsymbol{x}+t:=(x_{1}+t,\dots,x_{n}+t)$; see [3, Chapter 5.6]. This property holds, for example, if the kernel is shift-invariant, $K(x,y)=K(x+t,y+t)=:C(x-y)$, as in the case of the Poisson kernel, $K(x,y)=\mathbbm{1}_{\left\{x=y\right\}}$, or the Ginibre (Gaussian) kernel: $K(x,y)=\exp\left(-(x-y)^{2}/2\right)/\pi$. A result by Lavancier et al. [23, Theorem 1] states that $C$ can be any continuous correlation function with eigenvalues in its spectral representation bounded by 1, with examples provided in their work.
Theorem 4.5.
Assume that immigration is spanned by a stationary $(\Lambda,K)$-DPP with $\Lambda(\mathrm{d}x)=\lambda(x)\mathrm{d}x$, and that $\lambda(x)\sim\lambda_{\infty}e^{\delta x}$.
  1. 1.
    For all $\rho$, if $\delta>\max\{\rho,0\}$, then
    $\frac{{Z_{i}}(t)}{e^{\delta t}}\xrightarrow[t\to\infty]{L^{2}}A_{i}:=K_{\ast}\lambda_{\infty}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x,$
    where $K_{\ast}:=K(0,0)$.
  2. 2.
    If the process is supercritical, that is, $\rho>0$, and $\delta=\rho$, then
    $\frac{{Z_{i}}(t)}{te^{\delta t}}\xrightarrow[t\to\infty]{L^{2}}A_{i}^{\prime}:=K_{\ast}\lambda_{\infty}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i}.$
  3. 3.
    Assume that the process critical, i.e. $\rho=0$, $\lambda$ is bounded, $\delta=\rho$, and the covariance function $K(x,0)$ of the DPP tends to 00 as $x\to\infty$. Then
[…]
    where $Y\sim Gamma(\lambda_{\infty}\beta,Q)$ with
    $Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\left.\frac{\partial^{2}G_{\boldsymbol{\nu_{i}}}}{\partial x_{j}\partial x_{k}}\right|_{\boldsymbol{x}=\boldsymbol{1}}v_{i}u_{j}u_{k},\qquad\beta:=\frac{\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle}{Q},$
    and $\Gamma(\alpha,\beta)$ is the Gamma law with shape $\alpha$ and rate $\beta$, that is, its density is, for $x>0$,
    $f_{\alpha,\beta}(x)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-\beta x}.$
  4. 4.
    If the process is subcritical, that is, $\rho<0$, and $\delta=0$, then
[…]
    where $\boldsymbol{X}$ is a random variable whose Laplace transform is available in (5.15).
The theorem is proved in Section 5.3
Remark 4.6.
The result of item 3 in the last theorem in the case of Poisson immigration is established in [17, Theorem 8] and, under the stronger assumption that all moments of $I_{i}$ exist, in [35, Theorem 3]. After a close inspection, we have observed that the two in fact do not match. This is due to two errata: in the proof of [17, Theorem 8], there is a missed $r^{\ast}$ on the second line of page 221. The error in [35, Theorem 3] is due to an incorrect application to the result of Weiner [47]. In his work, the eigenvectors $\boldsymbol{u^{W}}$ and $\boldsymbol{v^{W}}$ are defined as, respectively, the right and left eigenvector of the matrix composed of $\mathbb{E}\left[(\nu_{i})_{j}\right]-\mathbbm{1}_{\left\{i=j\right\}}$, and not as the eigenvectors of the generator matrix $A$, see (4.4). The normalisation is $\langle\boldsymbol{u^{W}},\boldsymbol{1}\rangle=1$ and $\langle\boldsymbol{u^{W}},\boldsymbol{v^{W}}\rangle=1$. It is then direct that,
$u_{i}^{W}=u_{i},\quad v_{i}^{W}=\frac{1}{\sum_{i}(u_{i}v_{i}/\mu_{i})}\frac{v_{i}}{\mu_{i}},\quad\text{and}\quad\sum_{i}\mu_{i}u_{i}^{W}v_{i}^{W}=\frac{1}{\sum_{i}(u_{i}v_{i}/\mu_{i})}.$ (4.8)
The definition of $\boldsymbol{u}$ and $\boldsymbol{v}$ in this work is aligned with these in Sevastyanov [40] and Athreya and Ney [2].
5 Proofs
5.1 Proof of Theorem 4.3
Proof of Theorem 4.3.
We start with the generating function of the process, similarly to (4.3), the generating function of the multi-type process,
$G_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\sum_{\boldsymbol{n}\in\mathbb{N}^{d}}\mathbb{P}\left(\boldsymbol{Z}(t)=\boldsymbol{n}\right)\prod_{i=1}^{d}s_{i}^{n_{i}},$
is equal to
$1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\prod_{i=1}^{n}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\Lambda(\mathrm{d}x_{1})\dots\Lambda(\mathrm{d}x_{n}).$ (5.1)
To obtain the moments, we differentiate with respect to $s_{i}$ and evaluate at $\boldsymbol{s}=\boldsymbol{1}$. Note that $01-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})=$ 0. Therefore,
$\ \mathbb{E}\left[Z_{i}(t)\right]=\left.\frac{\partial}{\partial s_{i}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}$ $=\int_{(0,t]}K(x,x)\left.\frac{\partial}{\partial s_{i}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
$=\int_{(0,t]}K(x,x)\mathbb{E}\left[Z_{\times,i}(t-x)\right]\Lambda(\mathrm{d}x),$
as the derivatives of the terms in (5.1) for $n>1$ would have a multiple $1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{1})$ for some $i$, which is zero when evaluated at $\boldsymbol{s}=\boldsymbol{1}$. Similarly for the second moments, only the terms for $n\leq 2$ are relevant, so
$\mathbb{E}\left[Z_{i}(t)Z_{j}(t)\right]=\left.\frac{\partial}{\partial s_{i}\partial s_{j}}G_{\boldsymbol{Z}}(t,\boldsymbol{s})\right|_{\boldsymbol{s}=\boldsymbol{1}}=\int_{(0,t]}K(x,x)\left.\frac{\partial}{\partial s_{i}\partial s_{j}}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s})\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x)$
$\qquad\quad\,\,\,-\frac{1}{2!}\int_{(0,t]^{2}}\det(K(x_{i},x_{j}))_{1\leq i,j\leq 2}\left.\frac{\partial}{\partial s_{i}\partial s_{j}}\left(\prod_{i=1}^{2}\left(1-G_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\right)\right|_{\boldsymbol{s}=\boldsymbol{1}}\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2})$
$=\int_{(0,t]}K(x,x)\mathbb{E}\left[Z_{\times,i}(t-x)Z_{\times,j}(t-x)\right]\Lambda(\mathrm{d}x)$
$\quad\,\,+\frac{1}{2}\int_{(0,t]}\Big{(}K(x_{1},x_{1})K(x_{2},x_{2})-K^{2}(x_{1},x_{2})\Big{)}\Big{(}2\mathbb{E}\left[Z_{\times,i}(t-x_{1})Z_{\times,j}(t-x_{2})\Big{)}\right]\Lambda(\mathrm{d}x_{1})\Lambda(\mathrm{d}x_{2}).$
Using the covariance formula gives exactly the second item of Theorem 4.3, which concludes the proof. ∎
5.2 Proof of Theorem 4.4
Proof of Theorem 4.4.
We will work with the Laplace transform of $\boldsymbol{Z}$,
$\mathcal{L}_{\boldsymbol{Z}}(t,\boldsymbol{s}):=\mathbb{E}\left[\exp\left(\langle\boldsymbol{Z}(t),\boldsymbol{s}\rangle\right)\right],$
and use a similar technique to [35, Theorem 3]. To facilitate the formalism, define $\boldsymbol{Z}(t)=0$ for $t<0$. Using (4.5), we have that for positive $x$,
$\frac{\boldsymbol{Z}_{\times}(t-x)}{e^{\rho t}}\xrightarrow[t\to\infty]{a.s.}\boldsymbol{v}\cdot e^{-\rho x}W_{\times},$
which gives the pointwise convergence of the Laplace transforms, for each $x>0$,
$\mathcal{L}_{\boldsymbol{Z}_{\times}}\left((t-x),\boldsymbol{s}e^{-\rho t}\right):=\mathbb{E}\left[\exp\left(-\langle\boldsymbol{Z}(t-x),\boldsymbol{s}e^{-\rho t}\rangle\right)\right]\xrightarrow[t\to\infty]{}\mathcal{L}_{\boldsymbol{v}{W}_{\times}}\left(\boldsymbol{s}e^{-\rho x}\right).$ (5.2)
By (3.2) and Theorem 4.2, we get that
$\mathcal{L}_{\boldsymbol{Z}}\left(t,e^{-\rho t}\boldsymbol{s}\right)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\Lambda(\mathrm{d}\boldsymbol{x}),$ (5.3)
where for readability we introduced the functions, for a vector $\boldsymbol{x}\in\mathbb{R}^{n}$ with $\boldsymbol{x}=(x_{1},\dots,x_{n})$,
$D(\boldsymbol{x}):=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n},\quad\text{and}\quad\Lambda(\mathrm{d}\boldsymbol{x}):=\Lambda(\mathrm{d}x_{1})\cdots\Lambda(\mathrm{d}x_{n}).$
From (5.2), we have that for each $\boldsymbol{x}\in\mathbb{R}^{n}$,
$\mathbbm{1}_{\left\{\boldsymbol{x}\in(0,t]^{n}\right\}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\xrightarrow[t\to\infty]{}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{v}W_{\times}}(\boldsymbol{s}e^{-\rho x})\right).$ (5.4)
To use the last result under the integral sign in (5.3), we need to provide a bound with an integrable function. First, by Hadamard’s inequality, see e.g. [3, (15.A.1)],
$D(\boldsymbol{x})=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\leq\prod_{i=1}^{n}K(x_{i},x_{i}).$ (5.5)
Second, note that as we are working with Laplace transforms of positive random variables, we can assume that the vector $\boldsymbol{s}$ has positive entries. Let $C_{\boldsymbol{s}}:=\max_{j}s_{j}/\min v_{j}$ so that for all $j$, $0\leq s_{j}\leq C_{\boldsymbol{s}}v_{j}$. We next bound as in [35, Section 3.1],
$\begin{split}e^{-\rho t}\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]&=e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]\leq C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho(t-x)}\boldsymbol{Z}_{\times}(t-x),\boldsymbol{v}\rangle\right]\\ &=C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle e^{-\rho\cdot 0}\boldsymbol{Z}_{\times}(0),\boldsymbol{v}\rangle\right]=C_{\boldsymbol{s}}e^{-\rho x}\mathbb{E}\left[\langle\boldsymbol{I}_{0},\boldsymbol{v}\rangle\right]=:C_{\boldsymbol{s}}e^{-\rho x},\end{split}$ (5.6)
where we used the fact that $(\langle e^{-\rho t}\boldsymbol{Z}_{\times}(t),\boldsymbol{v}\rangle)_{t\geq 0}$ is a martingale, see [2, p.209]. Therefore,
$\left|1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x,\boldsymbol{s}e^{-\rho t})\right|$ $=\mathbb{E}\left[1-\exp\left(-e^{-\rho t}\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right)\right]\leq\mathbb{E}\left[e^{-\rho t}\langle\boldsymbol{Z}_{\times}(t-x),\boldsymbol{s}\rangle\right]$
$\leq C_{\boldsymbol{s}}e^{-\rho x},$
where in the first inequality we used that $1-e^{-x}\leq x$ for positive $x$. Thus we get
$\left|\mathbbm{1}_{\left\{\boldsymbol{x}\in(0,t]^{n}\right\}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\right|\leq C_{\boldsymbol{s}}^{n}\prod_{i=1}^{n}K(x_{i},x_{i})e^{-\rho x_{i}}.$ (5.7)
The last is indeed a $\Lambda(\mathrm{d}\boldsymbol{x})$-integrable because by assumption $\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)$ is finite, so
$\begin{split}&C_{\boldsymbol{s}}^{n}\int_{(0,\infty)^{n}}\prod_{i=1}^{n}\det(K(x_{i},x_{i}))e^{-\rho x_{i}}\Lambda(\mathrm{d}\boldsymbol{x})=\left(C_{\boldsymbol{s}}\int_{(0,\infty)}e^{-\rho x}K(x,x)\Lambda(\mathrm{d}x)\right)^{n}<\infty.\end{split}$ (5.8)
Therefore, we can apply the dominated convergence theorem in (5.4), which gives that, for each $n$,
$\int_{(0,t]^{n}}D(\boldsymbol{x})$ $\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s}e^{-\rho t})\right)\Lambda(\mathrm{d}\boldsymbol{x})$
[…]
Substituting in (5.3), again by dominated convergence and the bound (LABEL:eq:_bound_K), we obtain that
$\mathcal{L}_{\boldsymbol{Z}}\left(t,e^{-\rho t}\boldsymbol{s}\right)\xrightarrow[t\to\infty]{}1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{v}W_{\times}}(\boldsymbol{s}e^{-\rho x_{i}})\right)\Lambda(\mathrm{d}\boldsymbol{x}).$
We note that the sum in the last limit is absolutely convergent by (LABEL:eq:_bound_K). Letting $\Phi$ be the underlying DPP, we can rewrite the last as
$\mathcal{L}_{\boldsymbol{Z}}\left(t,e^{-\rho t}\boldsymbol{s}\right)\xrightarrow[t\to\infty]{}\mathcal{L}_{\Phi}\left(-\ln\left(\mathcal{L}_{\boldsymbol{v}W_{\times}}\left(\boldsymbol{s}e^{-\rho x}\right)\right)\right),$
which is exactly the Laplace transform of
$\int_{(0,\infty)}e^{-\rho x}\Phi^{\ast}(\mathrm{d}x),$
where $\Phi^{\ast}(\mathrm{d}x)$ is the random measure $\sum_{k}\beta_{k}\delta_{T_{i}}$, where $\beta_{k}$ are iid $\mathbb{R}^{d}$ random variables with distribution $\boldsymbol{v}W_{\times}$ and $T_{i}$ are the atoms of the determinantal point process $\Phi$, see [3, Example 2.2.30]. This concludes the proof of Theorem 4.4. ∎
5.3 Proof of Theorem 4.5
Proof of Theorem 4.5.
For the first two items we use techniques from the works of Yanev et al. [15, 16, 42, 44], for which we will need the following lemma.
Lemma 5.1 (Lemmas 2 and 3 from [42]).
Let $\lambda_{\infty}$ and $\psi_{\infty}$ be positive real constants.
  1. 1.
    If $\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}e^{\delta t}$ with $\delta\in\mathbb{R}$ and $\int_{0}^{\infty}e^{-\delta x}\psi(x)\,\mathrm{d}x$ finite, then
    $\int_{0}^{t}\lambda(t-u)\psi(u)\mathrm{d}u\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda(t)\int_{0}^{\infty}e^{-\delta x}\psi(x)\mathrm{d}x.$
  2. 2.
    If $\lambda(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}$ and $\psi(t)\stackrel{{\scriptstyle\infty}}{{\sim}}\psi_{\infty}$, then
    $\int_{0}^{t}\lambda(t-u)\psi(u)\mathrm{d}u\stackrel{{\scriptstyle\infty}}{{\sim}}\lambda_{\infty}\psi_{\infty}t.$
For items 3 and 4 we follow an approach similar to the one chosen by Rabehasaina and Woo in [35].
Let us start with the observation that the stationarity assumption (4.7), applied for $n=1$, implies that $K(x,x)$ is constant over $x$. Denote its value by $K_{\ast}$.
(1) We have by Theorem 4.3 that $\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}\int_{0}^{t}\mathbb{E}\left[Z_{\times,i}(t-x)\right]\lambda(x)\mathrm{d}x$. Furthermore, by classical results, see e.g. [40, p.151] or [2, p. 203],
$\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}a_{i}:=e^{\rho t}\langle\boldsymbol{u},\mathbb{E}\left[\boldsymbol{I}\right]\rangle v_{i}.$
Therefore, applying the first part of Lemma 5.1 with $\psi\equiv\mathbb{E}\left[Z_{\times,i}(\cdot)\right]$, we obtain
$\mathbb{E}\left[Z_{i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\delta t}A_{i}=:e^{\delta t}K_{\ast}\lambda_{\infty}a_{i}\int_{0}^{\infty}e^{-\delta x}\mathbb{E}\left[Z_{\times,i}(x)\right]\mathrm{d}x.$
We prove that $Z_{i}(t)$ concentrates around its mean by bounding its variance: by the second part of Theorem 4.3,
$Var(Z_{i}(t))\leq K_{\ast}\int_{0}^{t}\mathbb{E}\left[\left(Z_{\times,i}(t-x)\right)^{2}\right]\lambda(x)\mathrm{d}x.$
The asymptotics of the factorial moments of $Z_{\times,i}(t)$ are available in [40, p.152] or [2, p.204], giving us, for some positive constants $B_{j,i}$,
$\mathbb{E}\left[\left(Z_{\times,i}(t)\right)^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}\begin{cases}e^{\rho t}B_{1,i}&\text{if $\rho<0$},\\ tB_{2,i}&\text{if $\rho=0$},\\ e^{2\rho t}B_{3,i}&\text{if $\rho>0$}.\end{cases}$
Therefore, using again the first part of Lemma 5.1, and its second part for the case $2\rho=\delta$, we get that in the case where $\delta>\max\{0,\rho\}$,
$\quad\text{Var}(Z_{i}(t))=\begin{cases}O(e^{\delta t})&\text{if $\rho\leq 0$},\\ O(e^{\max\{2\rho,\delta\}t})&\text{if $\rho>0$ and $\delta\neq 2\rho$},\\ O(te^{\delta t})&\text{if $2\rho=\delta>0$}.\end{cases}$
Since we have already seen that $\mathbb{E}\left[Z_{i}(t)\right]\sim e^{\delta t}A_{i}$ and $\delta>\max\{0,\rho\}$ by assumption, we infer that always
$\mathbb{E}\left[\left(\frac{Z_{i}(t)}{\mathbb{E}\left[Z_{i}(t)\right]}-1\right)^{2}\right]=\frac{Var(Z_{i}(t))}{\mathbb{E}\left[Z_{i}(t)\right]^{2}}\xrightarrow[t\to\infty]{}0$ (5.9)
as claimed.
(2) Similarly to (1), Using the previously referred results of Sevastyanov, and Athreya and Ney, we have that, as $\delta=\rho>0$,
$\mathbb{E}\left[Z_{\times,i}(t)\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{\rho t}a_{i},\quad\text{and}\quad\mathbb{E}\left[(Z_{\times,i}(t))^{2}\right]\stackrel{{\scriptstyle\infty}}{{\sim}}e^{2\rho t}\left(B_{3,i}-a_{i}^{2}\right).$
Using again Lemma 5.1, we get that
$\mathbb{E}\left[Z_{i}(t)\right]=K_{\ast}\int_{0}^{t}\mathbb{E}\left[Z_{\times,i}(t-x)\right]\lambda(x)\mathrm{d}x\sim te^{\rho t}A^{\prime}_{i}:=te^{\rho t}K_{\ast}\lambda_{\infty}a_{i},\quad\text{and}\quad Var(Z_{i}(t))=\mathrm{O}\left(e^{2\rho t}\right).$
Therefore, the argument from (5.9) is applicable which establishes the result.
(4) In this case we assume that the process is subcritical ($\rho<0$) and that $\lambda(x)\sim\lambda_{\infty}$. By assumption
$\begin{split}\mathcal{L}_{\boldsymbol{Z}}\left(t,\boldsymbol{s}\right)&=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(t-x_{i},\boldsymbol{s})\right)\Lambda(\mathrm{d}\boldsymbol{x}).\end{split}$ (5.10)
To exploit the convergence of $\lambda(t)$ as $t\to\infty$, we change variables $x_{i}\to t-x_{i}$ in the last integral. Note that
$\det(K(t-x_{i},t-x_{j}))_{1\leq i,j\leq n}=\det(K(x_{i},x_{j}))_{1\leq i,j\leq n}\leq\prod_{i=1}^{n}K(x_{i},x_{i})=K_{\ast}^{n},$ (5.11)
where the first follows from our assumption that the DPP is stationary and also using inequality (5.5). Next, for $\boldsymbol{x}=(x_{1},\dots,x_{n})$, denoting
$\lambda(t-\boldsymbol{x}):=\lambda(t-x_{1})\dots\lambda(t-x_{n}),\quad\text{and}\quad\mathrm{d}\boldsymbol{x}:=\mathrm{d}x_{1}\dots\mathrm{d}x_{n},$
we get that
$\begin{split}\mathcal{L}_{\boldsymbol{Z}}\left(t,\boldsymbol{s}\right)&=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,t]^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\lambda(t-\boldsymbol{x})\mathrm{d}\boldsymbol{x}.\end{split}$ (5.12)
To apply the dominated convergence theorem, first bound as in (5.6) and with the same $C_{\boldsymbol{s}}$ as there,
$\left|1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x,\boldsymbol{s})\right|=\mathbb{E}\left[1-\exp\left(\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right)\right]\leq\mathbb{E}\left[\langle\boldsymbol{Z}_{\times}(x),\boldsymbol{s}\rangle\right]\leq e^{\rho x}C_{\boldsymbol{s}}.$ (5.13)
Next, pick $C$ such that $\sup_{t\geq C}\lambda(t)<\infty$, which exists as $\lambda$ has a finite limit at infinity. Therefore, for each $\boldsymbol{x}\in\mathbb{R}^{n}$,
$\begin{split}&\left|\mathbbm{1}_{\left\{\boldsymbol{x}\in(0,t]^{n}\right\}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\lambda(t-\boldsymbol{x})\right|\leq\prod_{i=1}^{n}\mathbbm{1}_{\left\{x_{i}\in(0,t)\right\}}K_{\ast}C_{\boldsymbol{s}}e^{\rho x_{i}}\lambda(t-x_{i})\\ &\hskip 28.45274pt=K_{\ast}^{n}C_{\boldsymbol{s}}^{n}\prod_{i=1}^{n}\left(\mathbbm{1}_{\left\{x_{i}\in(0,t-C)\right\}}e^{\rho x_{i}}\lambda(t-x_{i})+\mathbbm{1}_{\left\{x_{i}\in(t-C,t)\right\}}e^{\rho x_{i}}\lambda(t-x_{i})\right)\\ &\hskip 28.45274pt\leq K_{\ast}^{n}C_{\boldsymbol{s}}^{n}\prod_{i=1}^{n}\left(e^{\rho x_{i}}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x_{i})\right).\end{split}$
The last is an $\mathrm{d}\boldsymbol{x}$-integrable function with
$\begin{split}\int_{(0,\infty)^{n}}K_{\ast}^{n}C_{\boldsymbol{s}}^{n}&\prod_{i=1}^{n}\left(e^{\rho x_{i}}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x_{i})\right)\mathrm{d}\boldsymbol{x}\\ &=\left(K_{\ast}C_{\boldsymbol{s}}\int_{0}^{\infty}\left(e^{\rho x}\sup_{x\geq C}\lambda(x)+\mathbbm{1}_{\left\{x_{i}\in(0,C)\right\}}\lambda(x)\right)\mathrm{d}x\right)^{n}<\infty,\end{split}$ (5.14)
because by assumption the directing measure of a DPP $\Lambda$ is always $\sigma$-finite and because in this case $\rho<0$. Therefore by dominated convergence
$\int_{(0,t)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}$ $\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\lambda(t-\boldsymbol{x})\mathrm{d}\boldsymbol{x}$
$t\to\infty]{}\lambda_{\infty}^{n}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\mathrm{d}\boldsymbol{x},$
and again by dominated convergence, which we can apply thanks to the geometric bound (5.14), for the sum in (5.12), we get
$\mathcal{L}_{\boldsymbol{Z}}\left(t,\boldsymbol{s}\right)\xrightarrow[t\to\infty]{}1+\sum_{n\geq 1}\frac{(-1)^{n}\lambda^{n}_{\infty}}{n!}\int_{(0,\infty)^{n}}D(\boldsymbol{x})\prod_{i=1}^{n}\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}(x_{i},\boldsymbol{s})\right)\mathrm{d}\boldsymbol{x}.$ (5.15)
To conclude that the last is the Laplace transform of a positive random variable, it remains to verify that it is continuous at one, see e.g. [11, p. 431, Chapter XIII.1, Theorem 2]. This is indeed true by the bound (5.13) and the choice $C_{\boldsymbol{s}}=\max_{j}{s_{j}}/\min_{j}{v_{j}}$.
(3) For critical processes without immigration, the limit distribution is obtained under optimal assumptions (and for more general branching mechanisms) by Holte [14]. Again, see Remark 4.6, one should take into account the difference between the definition of $\boldsymbol{v}$ in [14] and in this work, notably that there are considered the eigenvectors of $\mathbb{E}\left[(\nu_{i})_{j}\right]$, compare with (4.4). Denoting with upper index $H$ the corresponding quantities from the work of Holte, and taking into account the normalization there $\langle\boldsymbol{u^{H}},\boldsymbol{v^{H}}\rangle=1,\langle\boldsymbol{v^{H}},\boldsymbol{1}\rangle=1$, we have that
$u_{i}^{H}=\frac{\sum_{j}(v_{j}/\mu_{j})}{\sum_{j}(u_{j}v_{j}/\mu_{j})}u_{i},\quad v_{i}^{H}=\frac{1}{\sum_{j}(v_{j}/\mu_{j})}\frac{v_{i}}{\mu_{i}},\quad\text{and}\quad\sum_{j}\mu_{j}u_{j}^{H}v_{j}^{H}=\frac{1}{\sum_{j}(u_{j}v_{j}/\mu_{j})}.$
Therefore from [14, Lemma on page p.493], we have that
$t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t,\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}{\beta}\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{Q+\boldsymbol{\langle s,\boldsymbol{v}\rangle}},$
where
$\beta:=\frac{\mathbb{E}\left[\langle\boldsymbol{I},\boldsymbol{u}\rangle\right]}{Q}\quad\text{with}\quad Q:=\frac{1}{2}\sum_{i,j,k=1}^{d}\frac{\partial^{2}G_{\boldsymbol{\nu_{i}}}}{\partial x_{j}\partial x_{k}}{v}_{i}{u}_{j}{u}_{k}.$
We will use the last limit in a form which is obtained by substituting, for $x\in(0,1)$ $t$ with $t(1-x)$, and $\boldsymbol{s}$ with $\boldsymbol{s}(1-x)$, which gives
$t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t(1-x),\frac{\boldsymbol{s}}{t}\right)\right)\xrightarrow[t\to\infty]{}\beta\frac{\langle\boldsymbol{s,\boldsymbol{v}}\rangle}{Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}.$
Substituting in (5.10) and changing variables $x_{i}\to tx_{i}$, we get
$\mathcal{L}_{\boldsymbol{Z}}\left(t,\frac{\boldsymbol{s}}{t}\right)=1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\int_{(0,1)^{n}}D(t\boldsymbol{x})\prod_{i=1}^{n}t\left(1-\mathcal{L}_{\boldsymbol{Z}_{\times}}\left(t(1-x_{i}),\frac{\boldsymbol{s}}{t}\right)\right)\lambda\left(t\boldsymbol{x}\right)\mathrm{d}\boldsymbol{x}$
Next, (5.11), (5.13) gives constant bounds of $D(t\boldsymbol{x})$ and the product. Moreover, as $\lambda$ is integrable, it is bounded for $||x||>\epsilon$. We also have that $D(t\boldsymbol{x})$ tends to 1 $\mathrm{d}\boldsymbol{x}$-a.e., because for $\boldsymbol{x}$ with different entries, $K(tx_{i},tx_{j})=K(t(x_{i}-x_{j}),0)$ tends to 0 if $x_{i}\neq x_{j}$ for $x\to\infty$ by assumption. Therefore by dominated convergence,
$\mathcal{L}_{\boldsymbol{Z}}\left(t,\frac{\boldsymbol{s}}{t}\right)$ $t\to\infty]{}1+\sum_{n\geq 1}\frac{(-1)^{n}}{n!}\left(\int_{0}^{1}\lambda_{\infty}\beta\frac{\langle\boldsymbol{s},\boldsymbol{v}\rangle}{Q+(1-x)\boldsymbol{\langle s,\boldsymbol{v}\rangle}}\mathrm{d}x\right)^{n}$
$=\left(\frac{Q}{Q+\langle\boldsymbol{s},\boldsymbol{v}\rangle}\right)^{\lambda_{\infty}\beta}.$
The last is indeed the Laplace transform of $Y\boldsymbol{v}$ with $Y\sim Gamma(\lambda_{\infty}\beta,Q)$ as claimed, which concludes the proof.
Acknowledgments
This study is financed by the European Union’s NextGenerationEU, through the National Recovery and Resilience Plan of the Republic of Bulgaria, project No BG-RRP-2.004-0008.